v_f= √(2 (-9.8 m/s^2 )(-25m))
1
Expert's answer
2020-05-11T20:11:28-0400

This is an equation that can be obtained from the law of conservation of energy: potential energy transforms into kinetic energy.


"mgh=\frac{1}{2}mv^2,\\
v=\sqrt{2gh}=v_f= \sqrt{2 \cdot9.8 \cdot25}=22\text{ m/s}."